find the x-intercept for the parabola defined by this equation y=4x^2-9x+5
Answer by lwsshak3(11628)

find the x-intercept for the parabola defined by this equation:
y=4x^2-9x+5
set y=0
4x^2-9x+5=0
(x-1)(4x-5)=0
x-intercepts
x=1
or
x=5/4